Prove the following result by contradiction:
Let f : X TO Y be a mapping. Suppose f (A INTERSECTION B) = f (A) INTERSECTION f (B) for all subsets
A, B PROPER SET X , f (PHI) = PHI. Then f is a 1-1 mapping.
1
Expert's answer
2020-05-21T14:47:00-0400

 Suppose it weren't necessary that "f" is injective. Then there would exist a non-injective function "f" such that for any two subsets "A,B\u2286X" we get "f(A\u2229B)=f(A)\u2229f(B)".

Let "a,b\in X"

"f(\{a\})\cap f(\{b\})=\{f(a)=f(b)\}\ne f(\{a\}\cap\{b\})\\
=f(\emptyset)=\emptyset"

Contradiction.


Suppose it weren't necessary that "f" is surjective.

Now assume "a \not\in f^{−1}(\{a\})". Then by definition of inverse image of set, "f(a) \not\in \{a\}" meaning "a\ne f(a)". Contradiction.


''injective + surjective = 1-1''